# Automated Multi-Task Learning for Joint Disease Prediction on Electronic Health Records

### v1

@@ Abstract.
In the realm of big data and digital healthcare, Electronic Health Records (EHR) have become a rich source of information with the potential to improve patient care and medical research. In recent years, machine learning models have proliferated for analyzing EHR data to predict patients’ future health conditions. Among them, some studies advocate for multi-task learning (MTL) to jointly predict multiple target diseases for improving the prediction performance over single task learning. Nevertheless, current MTL frameworks for EHR data have significant limitations due to their heavy reliance on human experts to identify task groups for joint training and design model architectures. To reduce human intervention and improve the framework design, we propose an automated approach named AutoDP, which can search for the optimal configuration of task grouping and architectures simultaneously. To tackle the vast joint search space encompassing task combinations and architectures, we employ surrogate model-based optimization, enabling us to efficiently discover the optimal solution. Experimental results on real-world EHR data demonstrate the efficacy of the proposed AutoDP framework. It achieves significant performance improvements over both hand-crafted and automated state-of-the-art methods, also maintains a feasible search cost at the same time [^1] Source code can be found via the link: https://github.com/SH-Src/AutoDP.
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456conference: KDD ’24: ACM SIGKDD International Conference on Knowledge Discovery & Data Mining; August 25-29, 2024; Barcelona, Spainbooktitle: KDD ’24: ACM SIGKDD International Conference on Knowledge Discovery & Data Mining, August 25-29, 2024, Barcelona, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Applied computing Health informatics
1. Introduction
In the era of big data and digital healthcare, the voluminous Electronic Health Records (EHR) have emerged as a treasure trove of valuable information that can revolutionize patient care, medical research, and clinical decision-making. As a result, the data mining community has been working on designing machine learning models to predict patients’ future health conditions by extracting patterns from their EHR data, such as mortality prediction (Ma et al., 2020), diagnosis prediction (Choi et al., 2016a; Ma et al., 2017) and hospital readmission (Huang et al., 2019).
@@ 1. Introduction
Although most of the existing machine learning based prediction models are designed to be single-task, i.e. predicting the risk of a single target disease, some of the existing works (Suo et al., 2017; Razavian et al., 2016; Wang et al., 2014; Zhao et al., 2022; Xu et al., 2019) proposed to design multi-task learning (MTL) models to jointly predict for multiple targets. The motivation behind this lies in the fact that two or more diseases might be related to each other in terms of sharing common comorbidities, symptoms, risk factors, etc. Consequently, training on related diseases simultaneously can offer additional insights and potentially enhance prediction performance. While multi-task learning offers potential advantages, the existing MTL frameworks for EHR data still suffer from the following limitations.
@@
Limitations of the existing MTL frameworks for EHR data To design an effective MTL framework, two fundamental challenges need to be addressed: (1) How can we determine which tasks should be trained together? (2) How can we design the model architecture for MTL?
@@
The first issue refers to the task grouping problem (Standley et al., 2020), which involves the selection of suitable task groups for joint training. Multi-task learning only provides advantages when the tasks are synergistic, i.e., training on the tasks together makes the model learn general knowledge that helps in performing the tasks better in the test set and prevents overfitting. Thus, given a large set of related tasks in a domain, we may need to group the tasks (allowing tasks to belong to multiple groups) together to create groups of tasks on each of which we will train a model. However, existing works usually rely on human expert discretion to select multiple tasks upfront and create a shared model for those tasks (Suo et al., 2017; Razavian et al., 2016; Wang et al., 2014; Zhao et al., 2022; Xu et al., 2019); hence, none of them has addressed the general problem of task grouping for EHR data. Moreover, due to the complexity of disease correlations, grouping synergistic tasks together is extremely challenging for human experts. It not only demands substantial effort (trying out every possible task combination) but also introduces the risk of task interference (putting disparate diseases together), potentially leading to performance degradation. Therefore, how to design the appropriate task grouping for MTL on EHR data presents a critical challenge.
@@
For the second concern, existing works typically rely on hand-crafted architectures for multi-task learning, which consist of a shared EHR encoder followed by several task-specific classifiers. However, due to the large number of possible operations as well as network topologies, it is near impossible to manually tune an optimal architecture for MTL. Furthermore, the optimal architectures for different task groups might also be distinct, so things can even get worse when the number of tasks grows and different task combinations are involved for joint tuning. Therefore, we need a more efficient and effective approach to design the optimal MTL architectures for EHR data.
@@
Automating the MTL framework design for EHR data To address the aforementioned challenges, we look to Automated Machine Learning (AutoML) (He et al., 2021) as a promising solution. Since AutoML relies on data-driven approaches to automate the design of machine learning algorithms, it has the potential to improve the design of MTL framework for EHR data by reducing human interventions. Moreover, several attempts have been explored in general domain (i.e. computer vision) for improving the design of task grouping (Fifty et al., 2021; Standley et al., 2020; Song et al., 2022) and MTL architectures (Ahn et al., 2019; Bragman et al., 2019; Sun et al., 2020; Guo et al., 2020; Gao et al., 2020). Therefore, AutoML has enormous promise to provide better solutions for MTL on EHR data. However, the exploration of AutoML in healthcare domain remains relatively limited (Waring et al., 2020). To the best of our knowledge, there are no existing work that automates the MTL framework design for EHR data, which is a notable gap in the field.
@@
Joint optimization over task grouping and architecture search More importantly, there is currently no end-to-end optimization framework for automating multi-task learning, even in the general domain. Current approaches separately address the problems of task grouping and architecture design. Firstly, a line of work (Fifty et al., 2021; Standley et al., 2020; Song et al., 2022) solves the task grouping problems by learning the task correlations. They operate under the underlying assumption that MTL architectures are the same across different task groups, which might not be practical. Secondly, researchers also apply Neural Architecture Search (NAS) (Elsken et al., 2019) to automatically design MTL architectures for a given set of tasks (Ahn et al., 2019; Bragman et al., 2019; Sun et al., 2020; Guo et al., 2020; Gao et al., 2020; Zhang et al., 2022), without considering the task grouping problem. However, optimizing these two aspects independently can lead to sub-optimal or even inaccurate estimations for the other, which, in turn, might lead to sub-optimal results for both aspects. Therefore, we need a more generalized AutoML framework for the joint optimization of both task grouping and architecture search. A significant challenge of this joint optimization is the enormous search space for all possible task combinations and candidate architectures. Brute force global optimization would demand significant computational resources and is infeasible. Hence, we need more efficient and effective solutions.
@@
Overview of the proposed approach Therefore, in this paper, we propose an Automated multi-task learning framework named AutoDP for joint Disease Prediction on electronic health records, which aims at jointly searching for the optimal task grouping and the corresponding neural architectures that maximize the multi-task performance gain. Specifically, we employ a surrogate model-based optimization approach (Liu et al., 2022) for efficient search. First, we define the joint search space of task combinations and architectures that includes all possible configurations for MTL. Our goal is to find the optimal solutions in this search space. To achieve that, the first question is how we can evaluate the performance of each configuration. Performing the ground truth evaluation for every configuration is infeasible, since it requires an entire multi-task learning procedure for each pair of architecture and task combination. Therefore, instead of exhaustively evaluating all the configurations, we build a surrogate model to predict the multi-task gains for any given configurations from the search space. In this way, we only need to evaluate the ground truth gains for a subset of samples from the search space, and use them to train the surrogate model for estimating the rest ones. The intuition is that there exists an underlying mapping from each configuration to the expected multi-task gains, thus it can be learned by a neural network. To train the surrogate model, we further propose an active sampling strategy to progressively collect training samples. This method is able to effectively train the surrogate model with only a small portion of ground truth evaluations from the search space. Eventually, after we obtain the trained surrogate model, we further use it to derive the final optimal task grouping and architectures. Because of the huge search space, it is not practical to use brute-force search. Therefore, we develop a greedy search method to find the near-optimal solution.
To sum up, our contributions are as follows:
@@
  • To the best of our knowledge, we are the first to propose an automated machine learning approach for multi-task learning on electronic health records AutoDP, which largely improves the design of task grouping and model architectures by reducing human interventions.
@@
  • We propose a surrogate model based optimization framework that is able to jointly search for the optimal task grouping and corresponding model architectures with high efficiency, which has not been solved by any previous works as well.
  • We propose an active sampling strategy to progressively construct the training set for the surrogate model, which improves sample efficiency by reducing the number of samples required for training the surrogate model.
@@
  • We develop a greedy search algorithm for deriving the final MTL configuration using the trained surrogate model, making it practical to find the near-optimal solution from the huge search space.
@@
  • Experimental results on real world EHR data - MIMIC IV (Johnson et al., 2020) demonstrate that the proposed AutoDP is capable of outperforming existing hand-crafted and automated methods under feasible computation cost.
2. Related Work
2.1. Multi-Task Learning with EHR
@@
Predictive modeling using electronic health records (EHR) (Ma et al., 2021) is a well-explored field. Researchers have applied deep learning methods to forecast patients’ future health conditions based on their historical EHR data. To enhance prediction performance further, some existing studies suggest employing multi-task learning framework to simultaneously predict multiple related target diseases or conditions, resulting in improved performance compared to single-task training. In particular, Wang, et al. (Wang et al., 2014) investigated the advantages of joint disease prediction using traditional machine learning models. More recently, researchers have applied recurrent neural network (RNN) based models to conduct multi-task learning on EHR data (Harutyunyan et al., 2019; Razavian et al., 2016; Suo et al., 2017), which is able to predict tasks like mortality, length of stay, ICD-9222https://www.cdc.gov/nchs/icd/icd9.htm diagnoses and etc. Additionally, Zhao, et al. (Zhao et al., 2022) also utilized a transformer based method for multi-task clinical risk prediction on multi-modal EHR data. However, all these studies manually select the set of tasks for joint training without task grouping and utilize a hand-crafted MTL model architecture, which largely limits their performance.
2.2. Multi-Task Grouping
@@
Multi-Task Grouping refers to the problem of identifying appropriate task groups for joint training. Due to the limitation of manually selected task groups, some of the work focus on obtaining the optimal task grouping through searching. Specifically, Standley, et al. (Standley et al., 2020) is the first work that systematically analyze the task correlations. For improving the efficiency, they use pair-wise MTL gains to estimate the high-order MTL gains, and obtain the pair-wise gains by training one model for each task pair. Based on the estimated gains, they derive the optimal task grouping using brute-force search. Fifty, et al. (Fifty et al., 2021) further improves the efficiency by training one model to derive all the pair-wise gains. They derive the task affinity based on the gradient information during training. Furthermore, Song, et al. (Song et al., 2022) propose a more general method that employs a meta model to learn the task correlations and estimates the high-order MTL gains more effectively. These works normally assume that the model architecture is the same across different task groups. But in practice, we can maximize performance gains by applying different model architectures with respect to each task group. Thus, we need a more general framework that considers the model architectures during task grouping.
2.3. Multi-Task NAS
@@
Neural Architecture Search (NAS) (Elsken et al., 2019) stands as a prominent research area in AutoML, focusing on the exploration of optimal deep network architectures through a data-driven approach. Popular search algorithm for NAS includes but is not limited to, reinforcement learning (Zoph and Le, 2016; Zoph et al., 2018), evolutionary search (Real et al., 2019), surrogate model based optimization (Liu et al., 2022) and differentiable methods (Liu et al., 2019). Although the main stream of NAS focuses on the setting of single task learning, some researchers also try to employ NAS in multi-task learning applications, predominantly for searching computer vision MTL architectures. Notably, studies done by Ahn, et al. (Ahn et al., 2019) and Bragman, et al. (Bragman et al., 2019) employ reinforcement learning and variational inference, respectively, to determine whether each filter in convolutions should be shared across tasks. Furthermore, other recent works (Sun et al., 2020; Guo et al., 2020; Gao et al., 2020; Zhang et al., 2022) leverage differentiable search algorithms (Liu et al., 2019), to determine the optimal sharing patterns across multiple network layers for diverse tasks. Despite the demonstrated advancements, a common limitation is their reliance on human experts to pre-define a set of tasks for joint training. This constraint poses challenges in practical scenarios where task grouping is not readily available, thereby limiting their broader applicability. What is more, their frameworks often search for better MTL architectures on top of one or several backbone architectures such as ResNet (He et al., 2016). However, such backbone architectures might not be available for EHR applications in medical domain. Therefore, a new multi-task NAS framework is needed for EHR data.
Figure 1. Overview of the proposed AutoDP
3. Methodology
3.1. Preliminaries
3.1.1. Problem definition
@@
Assume we have the input EHR data for multiple patients where each patient is represented as $\mathbf{X}\in\mathbb{R}^{L\times d_{e}}$, where $L$ is the time sequence length and $d_{e}$ is the hidden dimension of the input features. And for each patient, we have totally $N$ prediction tasks conditioned on the EHR data, denoted as $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{N}\}$. Our goal is to maximize the overall MTL performance gain for all these prediction tasks compared to single task training.
First, we make formal definition of MTL gain. We first conduct single task training on each task independently using a specific backbone model (such as RNN), and obtain the single-task performance for all tasks in terms of a predefined metric (such as averaged precision), denoted as $\{s_{1},s_{2},\cdots,s_{N}\}$. Then, the MTL gain is defined as:
(1) $g_{i}=\frac{(m_{i}-s_{i})}{s_{i}},i=1,\cdots,N,$
@@
In order to achieve that, our proposed method solves two searching problems at the same time using AutoML. First, we search for a list of task combinations that defines which tasks should be trained together. Second, we determine the optimal model architecture for each task combination. We aim at finding the optimal configuration for both, such that the highest overall gain $G$ is attained.
3.1.2. Task grouping search space
Given $N$ tasks, we can have totally $12^{N}-$ 1 task combinations, denoted as $\mathcal{C}=\{C_{1},C_{2},\cdots,C_{2^{N}-1}\}$, where every $C$ is a subset of $\mathcal{T}$. Given a budget $B$, we aim at searching for maximally $B$ task combinations from $\mathcal{C}$ to determine which tasks should be trained together such that the task combinations should cover all $N$ tasks so that we are able to obtain $\{m_{1},m_{2},\cdots,m_{N}\}$. If one task $T_{n}$ appears in multiple task combinations, we simply choose the highest performance for it as $m_{n}$.
3.1.3. Architecture search space
For every task combination, we also need to search for an MTL architecture to model the EHR data. We adopt the hard sharing mechanism as in most existing works (Harutyunyan et al., 2019; Suo et al., 2017), which consists of a shared encoder for extracting the latent representation of the input EHR and multiple task specific classifiers to generate the output for every task.
Specifically, we enable the search for the optimal shared encoder. For the search space of the encoder, we adopt the setting of directed acyclic graph (DAG) (Liu et al., 2019). The architecture is represented as a DAG that consists of $P$ ordered computation nodes, and each node is a latent feature that has connections to all previous nodes. For each connection (also called edge), we can choose one operation from a predefined set of candidate operations $\mathcal{O}$ for feature transformation. Let $\mathbf{E}^{0}=\mathbf{X}$, the formulation of node $p$ is defined as follows:
(2) $\mathbf{E}^{p}=\sum_{i=0}^{p-1}o_{(i,p)}(\mathbf{E}^{i}),o_{(i,p)}\in\mathcal{O},$
@@ 3.1.3. Architecture search space
To make final prediction, we take the last node representation $\mathbf{E}^{P}$ as the encoded feature for the input EHR, and use task specific classifiers to output final predictions, which are all fixed fully connected network layers.
3.1.4. MTL procedure
To evaluate a specific sample from the joint search space $\mathcal{C}\times\mathcal{A}$, we need to conduct an MTL experiment to obtain the multi-task performances. Specifically, given an architecture $A\in\mathcal{A}$ and a task combination $C\in\mathcal{C}$, we train the model $A$ to predict for all tasks in $C$ and get the multi-task performances for those tasks. Then, we can compute their gains by Eq.(1). In this way, we are able to evaluate how much gains that this sample $(C,A)$ could achieve.
3.2. Overview

### v2

@@ Abstract
Electronic Health Records (EHR) have become a rich source of information with the potential to improve patient care and medical research. In recent years, machine learning models have proliferated for analyzing EHR data to predict patients’ future health conditions. Among them, some studies advocate for multi-task learning (MTL) to jointly predict multiple target diseases for improving the prediction performance over single task learning. Nevertheless, current MTL frameworks for EHR data have significant limitations due to their heavy reliance on human experts to identify task groups for joint training and design model architectures. To reduce human intervention and improve the framework design, we propose an automated approach named AutoDP, which can search for the optimal configuration of task grouping and architectures simultaneously. To tackle the vast joint search space encompassing task combinations and architectures, we employ surrogate model-based optimization, enabling us to efficiently discover the optimal solution. Experimental results on real-world EHR data demonstrate the efficacy of the proposed AutoDP framework. It achieves significant performance improvements over both hand-crafted and automated state-of-the-art methods, also maintains a feasible search cost at the same time.
1 Introduction
@@
In the era of big data and digital healthcare, the voluminous Electronic Health Records (EHR) can revolutionize patient care, medical research, and clinical decision-making. Using these, the machine learning (ML) community has been designing models to predict patients’ future health conditions, e.g., models for mortality prediction [1], diagnosis prediction [2, 3] and hospital readmission [4]. Although most existing machine learning based prediction models are designed to be single-task, i.e. predicting the risk of a single target disease, some works [5, 6, 7, 8, 9] designed multi-task learning (MTL) models to jointly predict multiple targets. The motivation lies in the fact that two or more diseases might be related to each other in terms of sharing common comorbidities, symptoms, risk factors, etc. Consequently, training on related diseases simultaneously can offer additional insights and potentially enhance prediction performance. While multi-task learning offers potential advantages, the existing MTL frameworks for EHR data still suffer from the following limitations.
@@
Limitations of the existing MTL frameworks for EHR data. To design an effective MTL framework, two fundamental challenges need to be addressed:
@@
(1) How can we determine which tasks should be trained together? The task grouping problem [10] involves finding groups of tasks to train jointly. Multi-task learning only provides advantages when the tasks are synergistic, i.e., training on the tasks together makes the model learn general knowledge that helps in performing the tasks better in the test set and prevents overfitting. Thus, given a large set of related tasks in a domain, we may need to group the tasks (allowing tasks to belong to multiple groups) together to create groups of tasks on each of which we will train a model. However, existing works usually rely on human expert discretion to select multiple tasks upfront and create a shared model for those tasks [5, 6, 7, 8, 9]. Hence, none of them has addressed the general problem of task grouping for EHR data. Moreover, due to the complexity of disease correlations, grouping synergistic tasks together is extremely challenging for human experts. It not only demands substantial effort (trying out every possible task combination) but also introduces the risk of task interference (putting disparate diseases together), potentially leading to performance degradation. Therefore, how to design the appropriate task grouping for MTL on EHR data presents a critical challenge.
@@
(2) How can we design model architectures for MTL? Existing works [5, 6, 7, 8, 9] typically rely on hand-crafted architectures for multi-task learning, which consist of a shared EHR encoder followed by several task-specific classifiers. However, due to the large number of possible operations as well as network topologies, manually tuning an optimal architecture for MTL is impossible. Furthermore, the optimal architectures for different task groups might also be distinct. Thus, things can even get worse when the number of tasks grows and different task combinations are involved for joint tuning. Therefore, we need a more efficient and effective approach to design the optimal MTL architectures for EHR data.
@@
Automating the MTL framework design for EHR data. To address the aforementioned challenges, we look to Automated Machine Learning (AutoML) [11]. Since AutoML relies on data-driven approaches to automate the design of machine learning algorithms, it has the potential to improve the design of an MTL framework for EHR data and reduce human interventions. Several attempts have been explored in other domains, e.g., computer vision, to improve the design of task grouping [12, 10, 13] and MTL architectures [14, 15, 16, 17, 18]. However, the exploration of AutoML in healthcare domain remains relatively limited [19]. To the best of our knowledge, there are no existing work that automates the finding of groups of tasks for MTL towards designing an optimal framework for classification tasks using EHR data, which is a notable gap in the field.
@@
Joint optimization over task grouping and architecture search. Morever, currently there exists no end-to-end optimization framework for automating MTL, even in other domains. Current approaches independently address the problems of task grouping and architecture design. First, a line of work [12, 10, 13] solves the task grouping problems by learning the task correlations. They operate under the underlying assumption that MTL architectures are the same across different task groups, which might not be practical nor optimal. Second, researchers also apply Neural Architecture Search (NAS) [20] to automatically design MTL architectures for a predefined set of tasks [14, 15, 16, 17, 18, 21]. No existing work has integrated these two approaches to address both problems simultaneously. However, combining them naively could lead to sub-optimal results, as sequential optimization might result in inaccurate estimations for both aspects. Therefore, we need a more generalized AutoML framework for the joint optimization of both task grouping and architecture search.
Overview of the proposed approach. Therefore, in this paper, we show that an integrated approach for multi-task grouping and neural architecture search provides significant improvements. First, we extend existing single-task models like Retain [22], Adacare [16] to MTL in an EHR setting. Second, we apply DARTS [23], an NAS method used in MTL settings in different domains to the EHR domain. We use one shared model for predicting multiple tasks. These adaptations improve over the single-task setting. Second, we explore the impact of automated task grouping in the EHR setting by grouping tasks and finding an optimal NAS model for each task group. This further improves the performance. Finally, we propose an integrated framework an Automated multi-task learning framework, AutoDP, for joint Disease Prediction on electronic health records, which aims at jointly searching for the optimal task grouping and the corresponding neural architectures that maximize the multi-task performance gain. We show that this third method provides the maximal performance gain.
@@
Specifically, in AutoDP, we employ a surrogate model-based optimization approach [24] for efficient search. First, we define the joint search space of task combinations and architectures that includes all possible configurations for MTL. We want to find optimal solutions from this search space. To achieve that, the first question is how we can evaluate the performance of each configuration. Performing the ground truth evaluation for every configuration is infeasible, since it requires an entire multi-task learning procedure for each pair of architecture and task combination. Therefore, instead of exhaustively evaluating all the configurations, we build a surrogate model to predict the multi-task gains for any given configurations from the search space. In this way, we only need to evaluate the ground truth gains for a subset of samples from the search space, and use them to train the surrogate model for estimating the rest ones. The intuition is that there exists an underlying mapping from each configuration to the expected multi-task gains; thus it can be learned by a neural network. The remaining question is how we can effectively train the surrogate model using as few samples as possible. To this end, we further propose a progressive sampling strategy to guide the surrogate model training for improving sample efficiency. That is we train the surrogate model through multiple iterations. At every iteration, we select some points from the search space and update the surrogate model accordingly. The selection is conditioned on the current surrogate model and involves both exploitation and exploration. That is, we iteratively select the points that bring higher performance gains and also come from unexplored areas, which makes the training samples represent the whole search space. Eventually, after we obtain the trained surrogate model, we further use it to derive the final optimal task grouping and architectures. Because of the huge search space, it is not practical to use brute-force search. Hence, we develop a greedy search method to find the near-optimal solution.
In summary, our contributions are as follows:
@@
  • We are the first to propose an automated approach for multi-task learning on electronic health records AutoDP, which largely improves the design of task grouping and model architectures by reducing human interventions. Specifically, this work is the first to automate the design for the optimal task grouping and model architectures for MTL on EHR data.
@@
  • We are the first to propose a surrogate model based optimization framework that jointly searches for the optimal task grouping and corresponding model architectures with high efficiency in any domain.
  • We propose a progressive sampling strategy to construct the training set for the surrogate model, which improves sample efficiency by reducing the required number of ground truth evaluations during searching. Importantly, we balance exploitation and exploration so that the sampled configurations can represent the whole search space and are highly accurate.
@@
  • We propose a greedy search algorithm to derive the final MTL configuration using the trained surrogate model and find a near-optimal solution from the huge search space efficiently.
@@
  • Experimental results on real world EHR data - MIMIC IV [25] demonstrate that AutoDP improves classification performance significantly over existing hand-crafted and automated methods under feasible computational costs.
2 Related Work
@@
Multi-Task Learning with EHR. To enhance prediction performance while forecasting patients’ health conditions based on their historical data [26], existing studies employ multi-task learning to simultaneously predict multiple related target diseases or conditions, resulting in improved performance compared to single-task training. For example, Wang, et al. [7] investigated the advantages of joint disease prediction using traditional machine learning models. More recently, researchers have applied recurrent neural network (RNN) based models to conduct multi-task learning on EHR data [27, 6, 5], which is able to predict tasks like mortality, length of stay, ICD-9111https://www.cdc.gov/nchs/icd/icd9.htm diagnoses and etc. Additionally, Zhao, et al. [8] also utilized a transformer based method for multi-task clinical risk prediction on multi-modal EHR data. However, all these studies manually select the set of tasks for joint training without task grouping and utilize a hand-crafted MTL model architecture, which largely limits their performance.
@@
Multi-Task Grouping. Due to the limitation of manually selected task groups, some of the work focus on obtaining the optimal task grouping through searching. Specifically, Standley, et al. [10] is the first work that systematically analyze the task correlations. For improving the efficiency, they use pair-wise MTL gains to estimate the high-order MTL gains, and obtain the pair-wise gains by training one model for each task pair. Based on the estimated gains, they derive the optimal task grouping using brute-force search. Fifty, et al. [12] further improves the efficiency by training one model to derive all the pair-wise gains. They derive the task affinity based on the gradient information during training. Furthermore, Song, et al. [13] propose a more general method that employs a meta model to learn the task correlations and estimates the high-order MTL gains more effectively. These works normally assume that the model architecture is the same across different task groups. But in practice, we can maximize performance gains by applying different model architectures with respect to each task group. Thus, we need a more general framework that considers the model architectures during task grouping.
@@
Multi-Task NAS. Neural Architecture Search (NAS) [20] stands as a prominent research area in AutoML, focusing on the exploration of optimal deep network architectures through a data-driven approach. Although the main stream of NAS focuses on the setting of single task learning, some researchers also try to employ NAS in multi-task learning applications, predominantly for searching computer vision MTL architectures. Notably, studies done by Ahn, et al. [14] and Bragman, et al. [15] employ reinforcement learning and variational inference, respectively, to determine whether each filter in convolutions should be shared across tasks. Furthermore, other recent works [16, 17, 18, 21] leverage differentiable search algorithms [23], to determine the optimal sharing patterns across multiple network layers for diverse tasks. Despite the demonstrated advancements, a common limitation is their reliance on human experts to pre-define a set of tasks for joint training. This constraint poses challenges in practical scenarios where task grouping is not readily available, thereby limiting their broader applicability. What is more, their frameworks often search for better MTL architectures on top of one or several backbone architectures such as ResNet [28]. However, such backbone architectures might not be available for EHR applications in medical domain. Therefore, a new multi-task NAS framework is needed for EHR data.
3 Methodology
3.1 Preliminaries
@@
Problem definition. Assume we have the input EHR data for multiple patients where each patient is represented as $\mathbf{X}\in\mathbb{R}^{L\times d_{e}}$, where $L$ is the time sequence length and $d_{e}$ is the hidden dimension of the input features. We have $N$ prediction tasks using the EHR data, denoted as $\mathcal{T}=\{T_{1},T_{2},\cdots,T_{N}\}$. We seek to maximize the overall MTL performance gain for all these prediction tasks compared to single task training. First, we define MTL gain. Conduct a single task training on each task independently using a specific backbone model (such as RNN), and obtain the single-task performance for all tasks in terms of a predefined metric (such as average precision), denoted as $\{s_{1},s_{2},\cdots,s_{N}\}$. Then, the MTL gain is defined as:
$g_{i}=\frac{(m_{i}-s_{i})}{s_{i}},i=1,\cdots,N,$ (1)
@@
To achieve that, our proposed method solves two searching problems at the same time using AutoML. First, we search for a list of task combinations that defines which tasks should be trained together. Second, we determine the optimal model architecture for each task combination. We aim at finding the optimal configuration for both, such that the highest overall gain $G$ is attained.
Task grouping search space. For $N$ tasks, there are $12^{N}-$ 1 task combinations, $\mathcal{C}=\{C_{1},C_{2},\cdots,C_{2^{N}-1}\}$, where every $C$ is a subset of $\mathcal{T}$. Given a budget $B$, we aim at searching for maximally $B$ task combinations from $\mathcal{C}$ to determine which tasks should be trained together. The task combinations should cover all $N$ tasks so that we are able to obtain $\{m_{1},m_{2},\cdots,m_{N}\}$. If one task $T_{n}$ appears in multiple task combinations, we simply choose the highest performance for it as $m_{n}$.
Architecture search space. For every task combination, we also need to search for an MTL architecture to model the EHR data. We adopt the hard sharing mechanism as in most existing works [27, 5], which consists of a shared encoder for extracting the latent representation of the input EHR and multiple task specific classifiers to generate the output for every task.
Specifically, we enable the search for the optimal shared encoder. For the search space of the encoder, we adopt the setting of directed acyclic graph (DAG) [23]. The architecture is represented as a DAG that consists of $P$ ordered computation nodes, and each node is a latent feature that has connections to all previous nodes. For each connection (also called edge), we can choose one operation from a predefined set of candidate operations $\mathcal{O}$ for feature transformation. Let $\mathbf{E}^{0}=\mathbf{X}$, the formulation of node $p$ is defined as follows:
$\mathbf{E}^{p}=\sum_{i=0}^{p-1}o_{(i,p)}(\mathbf{E}^{i}),o_{(i,p)}\in\mathcal{O},$ (2)
@@
Finally, to predict, we take the last node representation $\mathbf{E}^{P}$ as the encoded feature for the input EHR, and use task-specific classifiers to output final predictions, which are all fixed fully connected network layers.
MTL procedure. To evaluate a specific sample from the joint search space $\mathcal{C}\times\mathcal{A}$, we need to conduct an MTL experiment to obtain the multi-task performances. Specifically, given an architecture $A\in\mathcal{A}$ and a task combination $C\in\mathcal{C}$, we train the model $A$ to predict for all tasks in $C$ and get the multi-task performances for those tasks. Then, we can compute their gains by Eq.(1). In this way, we are able to evaluate how much gains that this sample $(C,A)$ could achieve.
Figure 1: Overview of the proposed AutoDP
3.2 Overview